Blondie's Hair Extensions, Inc. makes hair extensions. They have provided the following data concerning its overhead costs for the coming year:

Wages and salaries $ 430,000
Depreciation 165,000
Rent 185,000
Total $ 780,000

Blondie's uses the activity-based costing system and have identified the activity pools "Assembly", "Order Processing" and "Other" and have provided the estimated level of activity for the upcoming year:

Activity Cost Pool Total Activity
Assembly 45,000 labor-hours
Order processing 600 orders
Other Not applicable

The "Other" activity cost pool does not have a measure of activity as it represents the accumulation of costs of idle capacity and organization-sustaining costs.

The distribution of resource consumption across activity cost pools is given below:

Activity Cost Pools
Assembly Order Processing Other Total
Wages and salaries 35% 30% 35% 100%
Depreciation 20% 50% 30% 100%
Rent 35% 30% 35% 100%

Given this information, the calculated activity rate for the Order Processing activity cost pool is closest to:

Multiple Choice

  • $435 per order

  • $353 per order

  • $455 per order

  • $445 per order
Answer #1
Order Processing
Wages and salaries 129000 =430000*30%
Depreciation 82500 =165000*50%
Rent 55500 =185000*30%
Total overhead allocated 267000
Total overhead allocated 267000
Divide by Total orders 600
Activity rate for the Order Processing activity 445
Option D $445 per order is correct